OpenStudy (anonymous):

Solve K = 1/3 (b – 2a) for b.

OpenStudy (btaylor):

3K = b-2a b=3K+2a

OpenStudy (anonymous):

thats the answer?

OpenStudy (btaylor):

yep

OpenStudy (anonymous):

thnk uuu

OpenStudy (anonymous):

first let us expand the right side so now we have K = (1/3)b -(2/3)a and this is simply (1/3)b = K +(2/3)a and now let us multiply by 3 to get b=3K +2a